OpenStudy (anonymous):

What is the distance between points (12, 4) and (18, –1)? Round to the nearest tenth.

OpenStudy (anonymous):

(18-12)^2+(-1-4)^2=x^2 x^2=36+25=61 x=sqrt61

OpenStudy (gorv):

((18-12)^2+(-1-4)^2)^1/2